# What If We Allocate Test-Time Compute Adaptively?

### v1

@@ 1 Introduction
Second, inference strategies are typically fixed a priori, for example, by pre-specifying the number of samples in Best-of-$N$ or the depth of a search procedure, rather than adapting based on intermediate reasoning quality (Rakhsha et al., 2025). As a result, the same exploration budget is applied regardless of whether early reasoning steps are clearly correct or already flawed.
@@ 2 Methodology
Problem Formulation. Given a problem $x$, the agent allocates inference-time computation by generating multiple reasoning trajectories across iterations. In each iteration, a specific configuration of reasoning tools (e.g., Chain-of-Thought, self-reflection, numeric or logical verification) and a compute strategy (e.g., Best-of-$N$, beam search, or lookahead with an exploration parameter) is selected, defining how the trajectory is generated. During generation, a PRM-based verifier evaluates intermediate reasoning steps and completed trajectories, assigning higher scores to locally consistent and mathematically valid transitions. These step-level signals are aggregated to prioritize promising reasoning paths, prune low-quality continuations, and ultimately favor trajectories that maintain consistent local correctness, which empirically leads to more reliable final answers. Each iteration yields a candidate trajectory, and the final output is selected via PRM-guided trajectory selection. Although we describe the framework using a trajectory-level decision formalism, the controller is not a learned policy. Instead, adaptivity is implemented entirely at inference time via deterministic, role-conditioned prompt templates applied to the same base LLM for planning, tool selection, compute strategy selection, verification, and answer extraction, without any pretraining, finetuning, or parameter updates as detailed in Appendix B. Finally, we analyze the compute metrics $F_{\text{theo}}$ and $S_{\text{CI}}$ to show that additional computation is selectively concentrated on high-utility reasoning trajectories rather than being uniformly scaled.
@@
3) Third, the compute selection agent $\mathcal{A}_{C}$ determines both the inference strategy $c$ and its exploration parameter $m\in\mathbb{N}^{+}$ via $c,m=\mathcal{A}_{C}(x,\pi,\mathcal{T}).$ Here, $m$ specifies the exploration budget associated with the chosen strategy: $N=m$ for Best-of-$N$ (BoN), beam width $k=m$ for Beam Search (BS), and lookahead depth $d=m$ for Lookahead Search (LA). The compute selector uses a structured prompt conditioned on $(x,\pi,\mathcal{T})$ to jointly select $c,m$ as detailed in Appendix B.4. These strategies trade off between exploration and exploitation: small $m$ commits early to a single trajectory, while larger $m$ explores multiple candidates to mitigate early errors at a higher compute cost. Jointly selecting $(c,m)$ enables per-problem adaptation, avoiding under-exploration on hard problems and wasted compute on easy ones. Tables 2 and 7 present comprehensive ablations across all tool-strategy-parameter combinations.
@@
For Best-of-$N$, we sample $N$ independent trajectories within iteration $\mathcal{I}_{i}$ and select $\hat{\tau}_{i}\;=\;\arg\max_{\tau\in\mathcal{S}_{i}}R(\tau).$ Here, the PRM is used to select the best completed trajectory within an iteration, rather than to prune partial prefixes during generation.
@@
where the sum runs over all models invoked at inference time, namely the base reasoning LLM, the LLM-based controller modules ($\mathcal{A}_{P}$, $\mathcal{A}_{T}$, $\mathcal{A}_{C}$), and the PRM. Here, $M_{j}$ denotes the parameter count of model $j$, $\bar{G}^{(j)}$ the average number of forward passes of model $j$ per problem, and $\bar{T}^{(j)}_{\text{total}}$ the average total tokens processed per forward pass, capped by the model context length $L^{(j)}_{\text{ctx}}$. For example, $\bar{G}^{(\text{base})}=N$ for Best-of-$N$ sampling, while $\bar{G}^{(\text{PRM})}$ counts all prefix- and trajectory-level PRM evaluations across iterations. This formulation ensures that raw compute from controller decisions and verification is fully included. While $F_{\text{theo}}$ captures overall computational scale, it does not distinguish effective reasoning from redundant work. To quantify compute utilization, we define the compute intensity score
@@ 4.2 Tool and Compute Strategy Selection Patterns
Figure 3 illustrates how the agent adapts reasoning tools and computation strategies under dynamic configuration. Tool usage shows clear model-dependent behavior: Llama spreads decisions across multiple tools (e.g., CoT and verifiers each used on the order of $\sim$10–15%), whereas Qwen concentrates heavily on structured reasoning plus numeric checking (e.g., CoT $\sim$40–50% and Numeric Verifier $\sim$30–40%). These patterns remain similar from 1 to 10 iterations, suggesting that tool selection is driven by model-specific reasoning characteristics rather than repeated sampling. Compute strategy usage is likewise complementary: Llama favors direct generation for a majority of instances ($\sim$60%), while Qwen allocates more mass to exploration strategies such as Best-of-$N$ and Lookahead (each often in the $\sim$15–35% range). Overall, these results indicate that dynamic configuration adapts inference behavior to model structure, steering each model toward reasoning paths that better align with its strengths rather than enforcing a single fixed inference recipe.
@@ 4.3 Comprehensive Ablation: Fixed Configurations
Compute efficiency. Dynamic + PRM Selection attains 81.4% accuracy while concentrating computation on high-utility trajectories and terminating weak ones early, resulting in substantially lower compute intensity despite slightly higher theoretical FLOPs. In contrast, as detailed in Appendix A.2, achieving the same accuracy with fixed configurations requires $1.31\times 10^{14}$ FLOPs and a compute intensity of $7.97\times 10^{-3}$ using CoT with Best-of-$N$ sampling at p=10p{=}10 and I=10I{=}10, reflecting significant wasted computation on easier instances.
Table 7 reports corresponding ablations on AIME24. Unlike MATH-500, no fixed configuration exceeds $10.0\%$ accuracy (best: CoT Lookahead, p=5p{=}5, $I\in\{1,5,10\}$), compared to $13.3\%$ for Dynamic + PRM Selection. Increasing $p$ degrades performance, for example, from $6.67\%$ to $10.0\%$ and then to $3.3\%$ for CoT Lookahead as $p$ increases, and gains from increasing $I$ are marginal. This indicates that trajectory diversity from adaptive multi-tool selection, rather than uniform repetition, is essential for highly challenging problems.
@@ 6 Related Work
Test-Time Compute Scaling. Scaling inference-time computation improves reasoning performance more effectively than increasing model size (Snell et al., 2024). Test-time compute scaling methods, including repeated sampling, beam search, chain-of-thought prompting, and verification-based selection, have substantially improved performance on reasoning-intensive tasks, particularly in mathematics (Wei et al., 2022; Lightman et al., 2023; Brown et al., 2024). Recent agentic approaches further integrate verification with structured search, such as Monte Carlo tree search with preference learning and self-refinement (Brown et al., 2024; Rakhsha et al., 2025; Xie et al., 2024; Zhang et al., 2024; Guan et al., 2025), but typically rely on fixed inference procedures (e.g., predetermined Best-of-$N$ or search depth), limiting flexibility across strategies. Complementary work explores adaptive allocation via difficulty estimation under fixed budgets (Du, 2025) or budget-aware tool invocation (Liu et al., 2025; Paglieri et al., 2025). More recent work studies adaptive inference policies that dynamically adjust reasoning depth, search breadth, or stopping criteria based on uncertainty or intermediate signals, rather than relying on fixed budgets or difficulty predictors (Banfi and Gamage, 2026).
@@ A.4 AIME24 Ablation Analysis
Selected Iteration Index: $1$ (out of $K=10$)

### v3

@@ 1 Introduction
Second, inference strategies are typically fixed a priori, for example, by pre-specifying the number of samples in Best-of-NN or the depth of a search procedure, rather than adapting based on intermediate reasoning quality (Rakhsha et al., 2025). As a result, the same exploration budget is applied regardless of whether early reasoning steps are clearly correct or already flawed.
@@ 2 Methodology
Problem Formulation. Given a problem $x$, the agent allocates inference-time computation by generating multiple reasoning trajectories across iterations. In each iteration, a specific configuration of reasoning tools (e.g., Chain-of-Thought, self-reflection, numeric or logical verification) and a compute strategy (e.g., Best-of-NN, beam search, or lookahead with an exploration parameter) is selected, defining how the trajectory is generated. During generation, a PRM-based verifier evaluates intermediate reasoning steps and completed trajectories, assigning higher scores to locally consistent and mathematically valid transitions. These step-level signals are aggregated to prioritize promising reasoning paths, prune low-quality continuations, and ultimately favor trajectories that maintain consistent local correctness, which empirically leads to more reliable final answers. Each iteration yields a candidate trajectory, and the final output is selected via PRM-guided trajectory selection. Although we describe the framework using a trajectory-level decision formalism, the controller is not a learned policy. Instead, adaptivity is implemented entirely at inference time via deterministic, role-conditioned prompt templates applied to the same base LLM for planning, tool selection, compute strategy selection, verification, and answer extraction, without any pretraining, finetuning, or parameter updates as detailed in Appendix B. Finally, we analyze the compute metrics $F_{\text{theo}}$ and $S_{\text{CI}}$ to show that additional computation is selectively concentrated on high-utility reasoning trajectories rather than being uniformly scaled.
@@
3) Third, the compute selection agent $\mathcal{A}_{C}$ determines both the inference strategy $c$ and its exploration parameter $m\in\mathbb{N}^{+}$ via $c,m=\mathcal{A}_{C}(x,\pi,\mathcal{T}).$ Here, $m$ specifies the exploration budget associated with the chosen strategy: $N=m$ for Best-of-NN (BoN), beam width $k=m$ for Beam Search (BS), and lookahead depth $d=m$ for Lookahead Search (LA). The compute selector uses a structured prompt conditioned on $(x,\pi,\mathcal{T})$ to jointly select $c,m$ as detailed in Appendix B.4. These strategies trade off between exploration and exploitation: small $m$ commits early to a single trajectory, while larger $m$ explores multiple candidates to mitigate early errors at a higher compute cost. Jointly selecting $(c,m)$ enables per-problem adaptation, avoiding under-exploration on hard problems and wasted compute on easy ones. Tables 2 and 7 present comprehensive ablations across all tool-strategy-parameter combinations.
@@
For Best-of-NN, we sample NN independent trajectories within iteration $\mathcal{I}_{i}$ and select $\hat{\tau}_{i}\;=\;\arg\max_{\tau\in\mathcal{S}_{i}}R(\tau).$ Here, the PRM is used to select the best completed trajectory within an iteration, rather than to prune partial prefixes during generation.
@@
where the sum runs over all models invoked at inference time, namely the base reasoning LLM, the LLM-based controller modules ($\mathcal{A}_{P}$, $\mathcal{A}_{T}$, $\mathcal{A}_{C}$), and the PRM. Here, $M_{j}$ denotes the parameter count of model $j$, $\bar{G}^{(j)}$ the average number of forward passes of model $j$ per problem, and $\bar{T}^{(j)}_{\text{total}}$ the average total tokens processed per forward pass, capped by the model context length $L^{(j)}_{\text{ctx}}$. For example, $\bar{G}^{(\text{base})}=N$ for Best-of-NN sampling, while $\bar{G}^{(\text{PRM})}$ counts all prefix- and trajectory-level PRM evaluations across iterations. This formulation ensures that raw compute from controller decisions and verification is fully included. While $F_{\text{theo}}$ captures overall computational scale, it does not distinguish effective reasoning from redundant work. To quantify compute utilization, we define the compute intensity score
@@ 4.2 Tool and Compute Strategy Selection Patterns
Figure 3 illustrates how the agent adapts reasoning tools and computation strategies under dynamic configuration. Tool usage shows clear model-dependent behavior: Llama spreads decisions across multiple tools (e.g., CoT and verifiers each used on the order of $\sim$10–15%), whereas Qwen concentrates heavily on structured reasoning plus numeric checking (e.g., CoT $\sim$40–50% and Numeric Verifier $\sim$30–40%). These patterns remain similar from 1 to 10 iterations, suggesting that tool selection is driven by model-specific reasoning characteristics rather than repeated sampling. Compute strategy usage is likewise complementary: Llama favors direct generation for a majority of instances ($\sim$60%), while Qwen allocates more mass to exploration strategies such as Best-of-NN and Lookahead (each often in the $\sim$15–35% range). Overall, these results indicate that dynamic configuration adapts inference behavior to model structure, steering each model toward reasoning paths that better align with its strengths rather than enforcing a single fixed inference recipe.
@@ 4.3 Comprehensive Ablation: Fixed Configurations
Compute efficiency. Dynamic + PRM Selection attains 81.4% accuracy while concentrating computation on high-utility trajectories and terminating weak ones early, resulting in substantially lower compute intensity despite slightly higher theoretical FLOPs. In contrast, as detailed in Appendix A.2, achieving the same accuracy with fixed configurations requires $1.31\times 10^{14}$ FLOPs and a compute intensity of $7.97\times 10^{-3}$ using CoT with Best-of-NN sampling at p=10p{=}10 and I=10I{=}10, reflecting significant wasted computation on easier instances.
Table 7 reports corresponding ablations on AIME24. Unlike MATH-500, no fixed configuration exceeds $10.0\%$ accuracy (best: CoT Lookahead, p=5p{=}5, $I\in\{1,5,10\}$), compared to $13.3\%$ for Dynamic + PRM Selection. Increasing $p$ degrades performance, for example, from $6.67\%$ to $10.0\%$ and then to $3.3\%$ for CoT Lookahead as $p$ increases, and gains from increasing II are marginal. This indicates that trajectory diversity from adaptive multi-tool selection, rather than uniform repetition, is essential for highly challenging problems.
@@ 6 Related Work
Test-Time Compute Scaling. Scaling inference-time computation improves reasoning performance more effectively than increasing model size (Snell et al., 2024). Test-time compute scaling methods, including repeated sampling, beam search, chain-of-thought prompting, and verification-based selection, have substantially improved performance on reasoning-intensive tasks, particularly in mathematics (Wei et al., 2022; Lightman et al., 2023; Brown et al., 2024). Recent agentic approaches further integrate verification with structured search, such as Monte Carlo tree search with preference learning and self-refinement (Brown et al., 2024; Rakhsha et al., 2025; Xie et al., 2024; Zhang et al., 2024; Guan et al., 2025), but typically rely on fixed inference procedures (e.g., predetermined Best-of-NN or search depth), limiting flexibility across strategies. Complementary work explores adaptive allocation via difficulty estimation under fixed budgets (Du, 2025) or budget-aware tool invocation (Liu et al., 2025; Paglieri et al., 2025). More recent work studies adaptive inference policies that dynamically adjust reasoning depth, search breadth, or stopping criteria based on uncertainty or intermediate signals, rather than relying on fixed budgets or difficulty predictors (Banfi and Gamage, 2026).
@@ A.4 AIME24 Ablation Analysis
Selected Iteration Index: 11 (out of $K=10$)